Find two consecutive integers such that the sum of 3 times the first integer and 8 times second integer is 107
Answer by checkley71(8403)

LET X & X=1 BE THE 2 COSECUTIVE INTEGERS.
3X+8(X+1)=107
3X+8X+8=107
11X=107-8
11X=99
X=99/11
X=9 ANSWER FOR THE FIRST INTEGER.
9+1=10 FOR THE SECOND INTEGER.
PROOF:
3*9+8*10=107
27+80=1-7
107=107